3. Find two different planes whose intersection is the line x = 1+t, y = 2-4,3 + 2t. Show work.
Answer #1

31 The ginen line is x = 1+t, y = 2-t and 2= 3+2t O) Set the two planes be Pi = Ax+By+GZ =D (2 Py = A x + B₂ y + G Z = D₂ ushx+2y + Ź Z = 13 2 he Putting the values of A1, B1, G. DI ; A2, B2, C2, Py in (2) we get P, = = 2x+4y + 2 =13 P2 = x+3y + 2 =